OpenStudy (anonymous):

find the difference (4y/xy)-(2x/xy^3)

OpenStudy (anonymous):

u must beging by simplify each expresion

OpenStudy (anonymous):

\[\frac{4y}{xy}-\frac{2x}{xy^3}\] \[=\frac{4y}{xy}\times \frac{y^2}{y^2}-\frac{2x}{xy^3}\] \[=\frac{4y^3}{xy^3}-\frac{2x}{xy^3}\] now that the denominators are the same, subtract in the numerator